y=x(2)-5 find the vertex tell if it opens up or down left or right and focus and directrix
Answer by khwang(438)

At least you should know how to type square correcly.
y=x^2-5,
or y+5 = x^2,
vertex is (0,-5).
Since y+5=x^2>=0, or y >= -5 the parabola opens up.
x^2 = 4*1/4(y+5)= 4a(y+5), a = 1/4
Focus = (0,-5+1/4) = (0,-19/4)
Directrix: y = -5 -1/4 = -21/4
Kenny